### v1

# Spherical Knot Mosaics

Abstract
In this paper we introduce the notion of a spherical knot mosaic where a knot is represented by tiling the surface of an $n\times n\times n$ topological $S^{2}$ with 11 canonical knot mosaic tiles and show this gives rise to several novel knot (and link) invariants: the spherical mosaic number, spherical tiling number, minimal spherical mosaic tile number, spherical face number, spherical $n$-mosaic face number, and minimal spherical mosaic face number. We show examples where this framework is an improvement over classical knot mosaics. Furthermore, we explore several bounds involving other classical knot invariants that these spherical mosaic invariants gives rise to.
KEYWORDS: knot, knot mosaic, spherical knot mosaic, spherical mosaic number, spherical tiling number, spherical face number
1 Introduction
Formally, a knot, $K$, is an equivalence class of a smooth embedding of $S^{1}$ into $S^{3}$ modulo the equivalence relation of smooth ambient isotopies. Given a representation of $K$, we can study $K$ by considering a two dimensional regular projection called a knot diagram where the only singularities allowed are transverse double crossings and each singularity denotes an over and under arc structure. These diagrams give rise to several knot invariants such as crossing number.
Knot mosaics impose further restrictions on these knot diagrams by having one represent a knot diagram using eleven $1\times 1$ mosaic tiles on an $n\times n$ grid. This additional structure allows one to define additional knot invariants such as the mosaic number. We discuss more of this background in Section 2.
In Section 3 we extend the notion of knot mosaics by introducing spherical knot mosaics where one considers a knot projected on $S^{2}\subset\mathbb{R}^{3}$ which we represent as the boundary of a cube with side length $n$ embedded as $0,n]\times[0,n]\times[0,n]\subset\mathbb{R}^{3}$ where each face is an $n\times n$ grid tiled by $1\times 1$ mosaic tiles and the edges of the cube connect adjacent faces. We then define the planar representation of a spherical knot mosaic which makes this an easier object to study and show examples of these for each of the first 35 unoriented prime knots in Appendix A.
In Section 4 we introduce several novel knot invariants using spherical knot mosaics and show examples where spherical knot mosaic invariants are stronger than classical knot mosaic invariants. In Section 5 we show several bounds on our spherical knot mosaic invariants involving classical knot invariants.
2 Background
A knot $n$-mosaic is a knot diagram formed on an $n\times n$ grid using 11 mosaic tiles depicted in Figure 1 with adjacent tiles suitably connected so that the black lines form a continuous path resulting in a knot diagram. The midpoint of the edge of a mosaic tile is a connection point if it is also the endpoint of a curve drawn on that tile. We call $T_{0}$ the empty tile and the other ten tiles non-empty tiles.
$T_{0}$ $T_{1}$ $T_{2}$ $T_{3}$ $T_{4}$ $T_{5}$ $T_{6}$ $T_{7}$ $T_{8}$ $T_{9}$ $T_{10}$
Figure 1: The 11 Mosaic Tiles
More formally, an $n$-mosaic is an $n\times n$ matrix such that each entry in the mosaic is one of the eleven mosaic tiles. When an $n$-mosaic has a single closed curve, then that curve is a knot diagram and the $n$-mosaic is called a knot $n$-mosaic [3]. A knot mosaic of a knot, $K$, can be denoted $M_{K}$.
In 2008, [5] first defined knot mosaics as a tool for representing quantum knots and conjectured that it could also be useful structure to study tame, classical knots. Later that year, [2] showed that mosaic knot theory and classical knot theory are equivalent. That is, it was shown that knot mosaic type is a complete knot invariant. Since that time knot mosaics have been widely used to study classical knots.
If an $n$-mosaic has greater than one closed component, then we call it a link $n$-mosaic. A link, $L$, is an equivalence class of smooth embedding of one or more disjoint copies of $S^{1}$ into $S^{3}$ modulo the equivalence relation of smooth ambient isotopies. All knots are also one component links. Links also give rise to link diagrams. Each component of a link diagram may either be linked with one or more other components or can be split into a separate components which has no crossings with any other component. For brevity’s sake we will only refer to knots throughout the rest of this paper, but many of the notions discussed can also be extended and defined for links.
Figure 2: Knot mosaics of the unknot (left), $4_{1}$ (center), and $7_{4}$ (right) [4]
In classical knot theory, one of the earliest and most widely studied knot invariants is the crossing number of a knot $K$, denoted $c(K)$, which is the minimum number of crossings among all possible knot diagrams representing $K$. Then through the lens of $n$-mosaics the most obvious invariant one can define with the structure of an $n$-mosaic is the mosaic number, $m(K)$, of a knot $K$. This is the smallest integer $n$ for which $K$ can be represented as a knot $n$-mosaic. For example, the unknot, $0_{1}$, has mosaic number equal to 2, meaning it cannot be represented on a 1-mosaic, but can on a 2-mosaic as in Figure 2. It is interesting to point out that for some knots it is not possible for a knot $n$-mosaic to achieve both $m(K)$ and $c(K)$ in the same diagram due to the restrictions imposed by the mosaic tiles.
One can also consider the knot tiling number, $t(K)$, which is the minimal number of non-empty tiles needed to represent knot a $K$ as a knot mosaic on a suitably sized mosaic. Closely related to these two knot invariants is the minimal mosaic tile number, $t_{M}(K)$, which is the minimum number of non-empty tiles required to construct a knot mosaic representing $K$ on a knot $m(K)$-mosaic.
3 Spherical Knot Mosaics
We extend the notion of a classical mosaic on an $n\times n$ grid by tiling the faces of an $n\times n\times n$ cube. Under this notion each face of the cube is an $n\times n$ grid we tile using the same eleven mosaic tiles. Since topologically, the cube is homeomorphic to $S^{2}$, we call such a tiling a spherical $n$-mosaic. When the tiling results in a single closed component, we call it a spherical knot $n$-mosaic. A spherical knot mosaic of a knot, $K$, can be denoted $SM_{K}$.
Figure 3: An example of the unknot as a spherical 1-mosaic using three tiles, $T_{2}$, $T_{3}$, and $T_{4}$.
Since it is both onerous and unenlightening to display more complicated 3D spherical knot mosaics in a 2D medium, we will break down each cube into a planar representation of $S^{2}$. This visualization will be formed by six $n\times n$ grids arranged in a $T$ shape with the appropriate edges identified.
Figure 4: Planar representation of the cube with identified edges color-coded.
We now have a variety of options for representing the unknot as a spherical knot 1-mosaic.
Figure 5: Three examples of a planar representation of the unknot.
Examples of planar representations of the first 35 non-trivial knots with crossing number less than or equal to 8 can be found in Appendix A.
4 Spherical Knot Mosaic Invariants
We now introduce several novel knot invariants using spherical knot mosaics. Our first novel knot invariant coming from spherical knot mosaics is an extension of the mosaic number. We define the spherical mosaic number of a knot, $K$, denoted $sm(K)$, to be the smallest integer, $n$, for which $K$ can be represented as a spherical knot $n$-mosaic. We can see in Appendix A that the spherical mosaic number for the trefoil and figure-eight knot is 1 and the $sm(K)\leq 2$ for all knots with $c(K)\leq 8$. We establish that the unknot, trefoil, and figure-eight knots are the only knots with spherical mosaic number 1 in Corollary 5.6.2.
This shows the the spherical knot mosaic number distinguishes $4_{1}$ and $5_{1}$ since $sm(4_{1})=1$ and $sm(5_{1})=2$. The mosaic number does not distinguish these knots as it is known that $m(4_{1})=m(5_{1})=5$ [4].
Our second novel invariant coming from spherical knot mosaics is an extension of the mosaic tiling number. Define the spherical tiling number of a knot, $K$, denoted $st(K)$, to be the smallest integer, $t$, for which $K$ can be represented as a spherical knot $n$-mosaic with $t$ non-empty mosaic tiles. One should note that this invariant does not require that the size of the mosaic $n$ be equal to $sm(K)$ when $st(K)$ tiles are used. Requiring such a restriction yields a third novel invariant based on the minimal mosaic tile number. We define the minimal spherical mosaic tile number of a knot, $K$, denoted $st_{M}(K)$ to be the smallest integer, $t_{M}$, for which $K$ can be represented as a spherical knot $sm(K)$-mosaic with $t_{M}$ non-empty mosaic tiles.
Proposition 4.1.
For a knot, $K$, $st(K)\leq st_{M}(K)$
Proof.
This follows directly from the definition. ∎
It is known that for knot mosaics there are knots which have different tiling numbers and minimal mosaic tile numbers [1]. It is quite possible this is also the case for spherical knot mosaics, however finding such an example may prove to be harder than it was for classical mosaics given how efficiently spherical mosaics use space compared to classical mosaics. This question is worth looking into in future research.
Three additional novel invariants which do not arise from classical knot mosaic invariants are the minimal number of faces of a cube required to represent a knot, the minimal number of faces of a cube to represent a knot for a fixed $n$ and the minimal number of faces required to represent a knot when $n=sm(K)$. Define the spherical face number of a knot, $K$, denoted $sf(K)$ to be the smallest integer, $f$, for which $K$ can be represented as a spherical knot $n$-mosaic with $f$ non-empty faces. This is a highly ineffective invariant given that it will be equal to 1 for all knots since every knot can be represented on an $n$-mosaic for $n\geq m(K)$ and thus fit on one face of a spherical $n$-mosaic for $n\geq m(K)$ as well.
Define the spherical $n$-mosaic face number of a knot, $K$, and fixed $n$, denoted $sf_{n}(K)$, to be the smallest integer, $f_{n}$, where $K$ can be represented as a spherical knot $n$-mosaic with $f_{n}$ non-empty faces.
Proposition 4.2.
For a knot $K$, if $n\geq m(K)$, then $sf_{n}(K)=1$.
Proof.
As each face of our spherical $n$-mosaic is at least an $m(K)$-mosaic by assumption, so there exists a planar mosaic representation of $K$ which can be glued to any $m(K)\times m(K)$ sub-mosaic of any face of our spherical $n$-mosaic. Thus, $sf_{n}(K)=1$. ∎
Remark 4.3.
Prop 4.2 means $sf_{n}(K)$ will only be a stronger invariant than $sf(K)$, and thus, nontrivial, when fixing an $n$ such that $m(K)>n\geq sm(K)$. We show that one can always choose such an $n$, that is, such an $n$ will exist that yields a valid spherical knot $n$-mosaic in Theorem 5.2 and Theorem 5.3.
Define the minimal spherical mosaic face number of a knot $K$, denoted $sf_{M}(K)$, to be the smallest integer, $f_{M}$, for which $K$ can be represented as a spherical knot $n$-mosaic with $f_{M}$ non-empty faces and $n=sm(K)$.
Example 4.4.
Lets see what values these invariants take for the unknot.
Knot
Invariant $0_{1}$
$sm(K)$ 1
$st(K)$ 3
$st_{M}(K)$ 3
$sf(K)$ 1
$sf_{1}(K)$ 3
$sf_{2}(K)$ 1
$sf_{M}(K)$ 3
Table 1: Some spherical mosaic theoretic invariants for $0_{1}$.
Proposition 4.5.
For any knot, $K$, $sf_{M}(K)=sf_{sm(K)}(K)$ and $1\leq sf_{n}(K)\leq sf_{M}(K)$ for all $n\geq sm(K)$.
Proof.
This follows directly from the definitions. ∎
This means that one can always represent a knot on a spherical $n$-mosaic for any $n$ larger than the minimum necessary the minimal number of faces required will decrease to 1 as $n$ increases. One could also consider a tuple made up of any number of these invariants. This will also be a knot invariant which may be stronger than any of these novel invariant alone.
5 Bounds on Spherical Knot Mosaic Invariants
In the above section we noted several bounds that naturally arise between our novel invariants. In this section we will explore bounds that arise between our invariants and preexisting knot invariants such as mosaic number, tiling number, and crossing number.
Proposition 5.1.
Given a knot, $K$, $sm(K)\leq m(K)$.
Proof.
Consider a knot, $K$, and its corresponding minimal $m(K)\times m(K)$ mosaic. Copy this mosaic onto any face of a spherical $m(K)$-mosaic. Thus, $m(K)$ is an upper bound for $sm(K)$. ∎
This first bound is obvious from the definition and can be improved upon. We next show that this bound is not sharp.
Theorem 5.2.
Given a knot, $K$, $sm(K)\leq m(K)-1$.
Proof.
Consider a knot, $K$, and its corresponding minimal $m(K)\times m(K)$ mosaic. Now consider any of the 4 possible $m(K)-1\times m(K)-1$ sub-mosaics. This will contain all but 2m(K)12m(K)-1 tiles of the original mosaic. Copy this sub-mosaic onto any face of a spherical (m(K)1)(m(K)-1)-mosaic and consider the remaining 2m(K)12m(K)-1 which are now hanging off two adjacent sides of our spherical (m(K)1)(m(K)-1)-mosaic. These two adjacent sides will have a shared corner tile. Delete it from the mosaic and glue the remaining 2m(K)22m(K)-2 non-corner boundary tiles to our spherical (m(K)1)(m(K)-1)-mosaic tiles on the appropriate adjacent cube faces.
The deleted tile will either be $T_{0}$ or one of $T_{1}$ through $T_{4}$. In either case, this operation of deleting the corner and mapping the remaining tiles to appropriate adjacent tiles on our cube preserves the knot type and closed-ness of our knot diagram. Therefore, we have successfully created a spherical knot (m(K)1)(m(K)-1)-mosaic of $K$ showing $sm(K)\leq m(K)-1$. ∎
Theorem 5.3.
Given a knot, $K$, $sm(K)\leq m(K)-2$.
Proof.
Let $K$ be a knot with mosaic number $m(K)$, and let $M_{K}$ be a planar mosaic representation of $K$ on an $m(K)\times m(K)$ grid. In a planar knot mosaic, all tiles with 4 connection points must be located within the interior $(m(K)-2)\times(m(K)-2)$ sub-mosaic to ensure suitable connectivity, so the remaining 4m(K)44m(K)-4 tiles on the boundary of our mosaic can be $T_{0}$ through $T_{6}$, but not any of $T_{7}$ through $T_{10}$.
Map the interior $(m(K)-2)\times(m(K)-2)$ sub-mosaic to any face of a spherical (m(K)2)(m(K)-2)-mosaic. There will be 4m(K)44m(K)-4 tiles hanging off the sides of our spherical (m(K)2)(m(K)-2)-mosaic.
Case 1: If a corner tile of $M_{K}$ is $T_{0}$, then we can remove it and preserve the knot type and closed-ness of our knot diagram by gluing the non-corner boundary tiles on the appropriate adjacent cube faces.
Case 2: If a corner of $M_{K}$ is one of $T_{1}$ through $T_{4}$, then we can still remove it and glue the non-corner boundary tiles on adjacent cube faces because doing so re-closes diagram as the two connection points previously touching the corner tile will now touch each other. Thus, omitting the corner tile and gluing the non-corner boundary tiles to the appropriate adjacent cube faces will still yield a continuous path preserving the knot type and closed-ness of our knot diagram.
This shows we have an upper bound on the spherical mosaic number in terms of the mosaic number:
$sm(K)\leq m(K)-2.$
The key idea in this proof suggests that further refinements of this bound might be possible depending on the spatial distribution of the crossings and the structure of the boundary tiles. For example, it may be that a planar mosaic representation of that knot where all the $T_{7}$, $T_{8}$, $T_{9}$, and $T_{10}$ tiles are located in a $k\times k$ sub-mosaic has $sm(K)\leq k$. However, this relies on the behavior of the paths between the connection points on the boundary of our sub-mosaic which while not having any crossings may complicate matters. This is a good direction for future research.
Proposition 5.4.
Given a knot, $K$, $st(K)\leq t(K)-1$.
Proof.
Consider an arbitrary knot $K$ and a corresponding planar $n$-mosaic representation which achieves the tiling number and has a $T_{3}$ tile such that there are no non-empty tiles in the entries both below or to the left of the $T_{3}$ tile. Such a representation will always exist but may not be a $m(K)$-mosaic. We can then represent $K$ on a spherical (n1)(n-1)-mosaic with t(K)1t(K)-1 tiles by deleting the tile with the $T_{3}$ tile, gluing all tiles above and to the right of the $T_{3}$ tile to the top face of the spherical (n1)(n-1)-mosaic such that the tile to upper right diagonal to the $T_{3}$ tile is glued to the bottom left corner of the top face of the spherical (n1)(n-1)-mosaic, gluing all tiles below or to the right of the $T_{3}$ tile to the front right face of the spherical (n1)(n-1)-mosaic such that the tile to the right of the $T_{3}$ tile is glued to the top left corner of the front right face of the spherical (n1)(n-1)-mosaic, and gluing all tiles above or to the left of the $T_{3}$ tile to the front left face of the spherical (n1)(n-1)-mosaic such that the tile above the $T_{3}$ tile is glued to the top right corner of the front left face of the spherical (n1)(n-1)-mosaic. ∎
Definition 5.5 (Spherical Mosaic Belt).
Let the boundary of a 3-cube with side length $n$ be tiled such that each face is an $n\times n$ mosaic, and all tiles conform to the standard mosaic tile set and connectivity rules. A spherical mosaic belt is a sequence of tiles satisfying the following conditions:
  1. 1.
    Each tile in the sequence is one of $T_{5},T_{6},T_{9}$ or $T_{10}$.
  2. 2.
    Consecutive tiles in the sequence are edge-adjacent and have connection points on opposite edges so that the sequence forms a continuous, non-branching closed path with no turns or gaps, but possibly under or over crossings.
  3. 3.
    Removing the tiles which form the mosaic belt disconnects the spherical mosaic.
Theorem 5.6.
If $SM_{K}$ is a spherical knot $n$-mosaic for a knot, $K$, then maximum number of $T_{9}$ or $T_{10}$ tiles of $SM_{K}$ is $6n^{2}-3n+1$.
Proof.
We will start with a spherical $n$-mosaic containing only crossing tiles, $T_{9}$ or $T_{10}$. There are $6n^{2}$ total tiles. This is not a spherical knot $n$-mosaic, but is a spherical link $n$-mosaic with exactly $3n$ connected components each forming a spherical mosaic belt. We will begin eliminating crossings one at a time until there is only one connected component.
Choose any spherical mosaic belt. Now choose any tile on that belt and replace it to a non-crossing tile with four connection points, $T_{7}$ or $T_{8}$. This change reduces the number of connected components in our spherical link $n$-mosaic by one and means the tile opposite the one we changed now represents a crossing of a single component instead of two disjoint components.
We can then iterate this process by choosing another crossing tile that is part of our larger connected component which does not represent the component crossing itself and changing it to a non-crossing tile with four connection points, this again reduces the number of connected components by one. We can do this until there is only one connected component and thus we have a spherical knot $n$-mosaic for some alternating knot $K$ with $6n^{2}-(3n-1)$ crossing tiles as there were originally $6n^{2}$ crossing tiles and we removed 3n13n-1 of them. This is the most crossings possible for a spherical knot $n$-mosaic. ∎
Corollary 5.6.1.
For each $n\in\mathbb{N}$ there is at least one knot, $K$, and representation of that knot as a spherical knot $n$-mosaic such that $K$ is alternating and
$6n^{2}-3n+1\geq c(K)$
Proof.
We modify the proof of Theorem 5.6 such that the original tiling with $6n^{2}$ crossing tiles alternates. That is, for each component every under crossing is preceded by and followed by an over crossing and vice versa. Now notice that replacing any crossing tile involving two components with $T_{7}$ or $T_{8}$ preserves the alternating nature of the link diagram, thus after replacing 3n13n-1 crossing tiles we are left with an alternating knot diagram with $6n^{2}-3n+1$ crossings. Note this may not be a reduced knot diagram so we do not necessarily find that this is the crossing number, but instead an upper bound for the crossing number of our knot. ∎
Figure 6: An example of a 22- spherical knot mosaic which has the maximal number of crossing tiles of $6\cdot 2^{2}-3\cdot 2+1=19$.
Corollary 5.6.2.
If $K$ is a knot with spherical mosaic number $sm(K)$, then $6(sm(K))^{2}-3(sm(K))+1\geq c(K)$.
From Corollary 5.6.2 we have that
Proposition 5.7.
Let $SM_{K}$ be a spherical knot $1$-mosaic for some knot $K$, then $4\geq c(K)$.
which further implies
Corollary 5.7.1.
Let $K$ be a knot such that $c(K)\geq 5$, then $sm(K)\geq 2$.
Similarly from Corollary 5.6.2 we have that
Proposition 5.8.
Let $SM_{K}$ be a spherical knot $2$-mosaic for some knot $K$, then $19\geq c(K)$.
which further implies
Corollary 5.8.1.
Let $K$ be a knot such that $c(K)\geq 20$, then $sm(K)\geq 3$.
There are several more questions spherical knot mosaics gives rise to which the authors plan to explore in subsequent papers.
Acknowledgments
The authors would like to thank Jennifer Schultens for her encouragement and advice throughout this project.
Appendix A Planar Representation Examples
Planar representations of the first 35 unoriented prime knots as spherical knot mosaics. Due to Corollary 5.6.2, each of these planar representations realize their knot type’s spherical mosaic number.
[Uncaptioned image][Uncaptioned image][Uncaptioned image]